# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 7.6. Proof of Part (e)
$\xi_{n-1-i}^{q^{i}}M(0,m)=\sum_{j=1}^{m}(-1)^{j+1}\xi_{n-1-i-j}^{q^{i}}M(j,m).$
From part (b), $u_{m}d_{j,m}=M(j,m)+\delta_{j,m}$ with $\nu(\delta_{j,m})>1+q+\cdots+q^{m-1}$. Substituting into the expression for $\gamma$ gives
$\gamma$ $=$ $\sum_{j=0}^{m-1}(-1)^{j}\xi_{n-2-i-j}^{q^{i}}M(j+1,m)+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+$ $\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}$
$=$ $\xi_{n-1-i}^{q^{i}}M(0,m)+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+$ $\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}.$
@@
$\gamma-\xi_{n-1-i}^{q^{i}}u_{m}$ $=$ $-\xi_{n-1-i}^{q^{i}}\delta_{0,m}+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+$ $\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i-1}}+\gamma_{n-2-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}.$
By induction, $\nu(\gamma_{j,m-1}^{(i-1)})>q^{i-1}$. Therefore
$\nu(\gamma-\xi_{n-1-i}^{q^{i}}u_{m})>q^{i}+1+q+\cdots+q^{m-1},$
giving $\nu(\gamma/u_{m}-\xi_{n-1-i}^{q^{i}})>q^{i}$. Define $\gamma_{n-1-i,m}^{(i)}:=\gamma/u_{m}-\xi_{n-1-i}^{q^{i}}$ and, for $m-1-i\leq\ell<n-1-i$, define $\gamma_{\ell,m}^{(i)}:=(\gamma_{\ell,m-1}^{(i-1)})^{q}$. Then
$\gamma/u_{m}=\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i}}+\gamma_{n-2-i-j,m}^{(i)})d_{j+1,m}$
@@
$\xi_{n-1-i}^{q^{i}}$ $=$ $\gamma/u_{m}-\gamma_{n-1-i,m}^{(i)}$
$=$ $\left(\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-2-i-j}^{q^{i}}+\gamma_{n-2-i-j,m}^{(i)})d_{j+1,m}\right)-\gamma_{n-1-i,m}^{(i)}$
$=$ $\left(\sum_{j=1}^{m}(-1)^{j+1}(\xi_{n-1-i-j}^{q^{i}}+\gamma_{n-1-i-j,m}^{(i)})d_{j,m}\right)-\gamma_{n-1-i,m}^{(i)}$
@@ 8. Generation Over The Steenrod Algebra
$S_{m}^{G_{m}}$ is generated as an $\mathcal{A}$-algebra by the two elements $\xi_{0}$ and $d_{1,m}$.
@@
Let $C$ denote the smallest algebra, closed under the action of the Steenrod algebra and containing $\{\xi_{0},d_{1,m}\}$. Clearly $C\subseteq S_{m}^{G_{m}}$. Since $\mathcal{P}^{q^{i}}(\xi_{i})=\xi_{i+1}$ we see that $\xi_{i}\in C$ for all $i\geq 0$. It remains to show that $d_{i,m}\in C$ for $i=1,2,3,\dots,m$. We proceed by induction. Suppose then that $d_{1,m},d_{2,m},\dots,d_{i-1,m}\in C$ and consider $F:=\mathcal{P}^{q^{m-i-1}}(d_{i-1,m})$. Recall, from the proof of Theorem 3.1, that $d_{i-1,m}\equiv_{I}\pm d_{i-1}(W_{1})$ where $I$ is the ideal $I=\langle x_{1},x_{2},\ldots,x_{m}\rangle$, $W_{1}={\rm Span}_{\mathbb{F}_{q}}\{y_{1},\ldots,y_{m}\}$, and $d_{i-1}(W_{1})$ is the $(i-1)^{\text{th}}$ Dickson invariant for $W_{1}$. Then $F\in S_{m}^{G_{m}}$ is homogeneous of degree $q^{m-1}-q^{m-i}$. Since $\mathcal{P}^{q^{m-i-1}}(d_{i-1}(W_{1}))=-d_{i}(W_{1})$ (see [24]) it follows that $F\equiv_{I}\mathcal{P}^{q^{m-i-1}}(d_{i-1}(W_{1}))=-d_{i}(W_{1})\neq 0$. But the only invariants of degree $q^{m-1}-q^{m-i}$ which do not vanish on $W_{1}$ are scalar multiples of $d_{i,m}$. Hence $F=\pm d_{i,m}+f$ where $f\equiv_{I}0$ and $f\in\mathbb{F}_{q}[\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{i-1,m}]\subseteq C$. This shows that $d_{i,m}\in C$ as required. ∎
@@
As in the proof of Theorem 3.11, let $I$ denote the ideal in $S_{m}$ generated by $\{x_{1},\ldots,x_{m}\}$. Let $J$ denote the ideal in $S_{m}^{G_{m}}$ generated by $\{\xi_{0},\ldots,\xi_{n-2}\}$. Since $J\subset I$, the inclusion of $S_{m}^{G_{m}}$ into $S_{m}$ induces a map from $S_{m}^{G_{m}}/J$ to $S_{m}/I$. Since both $I$ and $J$ are closed under the action of the Steenrod algebra, this is a map of $\mathcal{A}$-algebras. Let $D$ denote the subalgebra of $S_{m}$ generated by $\{d_{1,m},\ldots,d_{m,m}\}$, let $\widetilde{D}$ denote the image of $D$ in $S_{m}^{G_{m}}/J$ and let $\overline{D}$ denote the image of $D$ in $S_{m}/I$. As algebras $D$, $\widetilde{D}$, and $\overline{D}$ are isomorphic. Therefore the map from $S_{m}^{G_{m}}/J$ to $S_{m}/I$ restricts to an isomorphism $\widetilde{D}$ to $\overline{D}$ as $\mathcal{A}$-algebras. From the proof of Theorem 3.9, $\overline{D}$ is isomorphic to $(\mathbb{F}_{q}[W_{1}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}})^{q^{m-1}}$. Since $d_{1}(W_{1})$ generates $\mathbb{F}_{q}[W_{1}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}}$ (see [24]), $\overline{d_{1}}$ generates $\overline{D}$ and $\tilde{d_{1}}$ generates $\widetilde{D}$, as $\mathcal{A}$-algebras. From this we see that $\{\xi_{0},d_{1,m}\}$ generates $S_{m}^{G_{m}}$ as an $\mathcal{A}$-algebra. ∎
@@
$S_{m}^{G_{m}}$ is generated as an $\mathcal{A}$-algebra by $\{\xi_{0},\mathcal{R}(N(y_{1})^{q-1})\}$.
@@
Since $\deg(\mathcal{R}(N(y_{1})^{q-1}))=\deg(d_{1,m})<\deg(d_{i,m})$ for $i>1$, it follows from Theorem 4.6, that $\mathcal{R}(N(y_{1})^{q-1})=cd_{1,m}+\delta$ for some scalar $c$ and $\delta\in R_{n-2}$. By Theorem 5.21 and Lemma 3.10
@@ Acknowledgement
The computer algebra package Magma [2] was very useful in this work. It was used to test and confirm many hypotheses in low dimensions for small primes. We are grateful to Karl Dilcher and Keith Taylor (Canadian Mathematical Society book editors) and Donna Chernyk (Springer) for encouraging us to work on this problem.

### v3

@@ 7.4. Proof of Part (f)
$\xi_{n-i}^{q^{i}}M(0,m)=\sum_{j=1}^{m}(-1)^{j+1}\xi_{n-i-j}^{q^{i}}M(j,m).$
From part (d), $u_{m}d_{j,m}=M(j,m)+\delta_{j,m}$ with $\nu(\delta_{j,m})>1+q+\cdots+q^{m-1}$. Substituting into the expression for $\gamma^{(i)}$ gives
$\gamma^{(i)}$ $=$ $\sum_{j=0}^{m-1}(-1)^{j}\xi_{n-1-i-j}^{q^{i}}M(j+1,m)+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-1-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-1-i-j}^{q^{i-1}}+\gamma_{n-1-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}$
$=$ $\xi_{n-i}^{q^{i}}M(0,m)+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-1-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-1-i-j}^{q^{i-1}}+\gamma_{n-1-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}.$
@@
$\gamma^{(i)}-\xi_{n-i}^{q^{i}}u_{m}$ $=$ $-\xi_{n-i}^{q^{i}}\delta_{0,m}+\sum_{j=0}^{m-1}(-1)^{j}(\gamma_{n-1-i-j,m-1}^{(i-1)})^{q}M(j+1,m)$
$+\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-1-i-j}^{q^{i-1}}+\gamma_{n-1-i-j,m-1}^{(i-1)})^{q}\delta_{j+1,m}.$
By induction, $\nu(\gamma_{j+1,m-1}^{(i-1)})>q^{i-1}$. Therefore
$\nu(\gamma^{(i)}-\xi_{n-i}^{q^{i}}u_{m})>q^{i}+1+q+\cdots+q^{m-1},$
giving $\nu(\gamma^{(i)}/u_{m}-\xi_{n-i}^{q^{i}})>q^{i}$. Define $\gamma_{n-i,m}^{(i)}:=\gamma^{(i)}/u_{m}-\xi_{n-i}^{q^{i}}$ and, for $m-1-i\leq\ell<n-1-i$, define $\gamma_{\ell+1,m}^{(i)}:=(\gamma_{\ell+1,m-1}^{(i-1)})^{q}$. Then
$\gamma^{(i)}/u_{m}=\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-1-i-j}^{q^{i}}+\gamma_{n-1-i-j,m}^{(i)})d_{j+1,m}$
@@
$\xi_{n-i}^{q^{i}}$ $=$ $\gamma^{(i)}/u_{m}-\gamma_{n-i,m}^{(i)}$
$=$ $\left(\sum_{j=0}^{m-1}(-1)^{j}(\xi_{n-1-i-j}^{q^{i}}+\gamma_{n-1-i-j,m}^{(i)})d_{j+1,m}\right)-\gamma_{n-i,m}^{(i)}$
$=$ $\left(\sum_{j=1}^{m}(-1)^{j+1}(\xi_{n-i-j}^{q^{i}}+\gamma_{n-i-j,m}^{(i)})d_{j,m}\right)-\gamma_{n-i,m}^{(i)}$
@@ 8. Generation Over The Steenrod Algebra
$S_{m}^{G_{m}}$ is generated as an $\mathcal{A}$-algebra by the two elements $\xi_{1}$ and $d_{1,m}$.
@@
Let $C$ denote the smallest algebra, closed under the action of the Steenrod algebra and containing $\{\xi_{1},d_{1,m}\}$. Clearly $C\subseteq S_{m}^{G_{m}}$. Since $\mathcal{P}^{q^{i-1}}(\xi_{i})=\xi_{i+1}$ we see that $\xi_{i+1}\in C$ for all $i\geq 1$. It remains to show that $d_{i,m}\in C$ for $i=1,2,3,\dots,m$. We proceed by induction. Suppose then that $d_{1,m},d_{2,m},\dots,d_{i-1,m}\in C$ and consider $F:=\mathcal{P}^{q^{n-i-1}}(d_{i-1,m})$ of degree $q^{n-1}-q^{n-1-i}$. Recall, from the proof of Theorem 3.12 that $d_{i-1,m}\equiv_{I}\pm d_{i-1}(W_{m})^{q^{m-1}}$ where $I$ is the ideal $I=\langle x_{1},x_{2},\ldots,x_{m}\rangle$, $W_{m}={\rm Span}_{\mathbb{F}_{q}}\{y_{m},\ldots,y_{1}\}$, and $d_{i-1}(W_{m})$ is the $(i-1)^{\text{th}}$ Dickson invariant in the variables $y_{m},\ldots,y_{1}$. Since $\mathcal{P}^{q^{m-i}}(d_{i-1}(W_{m}))=-d_{i}(W_{m})$ (see [29]) it follows that $F\equiv_{I}\mathcal{P}^{q^{n-i-1}}(d_{i-1}(W_{1})^{q^{m-1}})=-d_{i}(W_{1})^{q^{m-1}}\neq 0$. But the only invariants of degree $q^{n-1}-q^{n-1-i}$ which do not vanish on $W_{m}$ are scalar multiples of $d_{i,m}$. Hence $F=\pm d_{i,m}+f$ where $f\equiv_{I}0$ and $f\in\mathbb{F}_{q}[\xi_{1},\ldots,\xi_{m},d_{1,m},\ldots,d_{i-1,m}]\subseteq C$. This shows that $d_{i,m}\in C$ as required. ∎
@@
As in the proof of Theorem 3.12, let $I$ denote the ideal in $S_{m}$ generated by $\{x_{1},\ldots,x_{m}\}$. Let $J$ denote the ideal in $S_{m}^{G_{m}}$ generated by $\{\xi_{1},\ldots,\xi_{n-1}\}$. Since $J\subset I$, the inclusion of $S_{m}^{G_{m}}$ into $S_{m}$ induces a map from $S_{m}^{G_{m}}/J$ to $S_{m}/I$. Since both $I$ and $J$ are closed under the action of the Steenrod algebra, this is a map of $\mathcal{A}$-algebras. Let $D$ denote the subalgebra of $S_{m}$ generated by $\{d_{1,m},\ldots,d_{m,m}\}$, let $\widetilde{D}$ denote the image of $D$ in $S_{m}^{G_{m}}/J$ and let $\overline{D}$ denote the image of $D$ in $S_{m}/I$. It follows from Theorem 3.12 and its proof that $D\cap J\subset D\cap I=0$. Therefore, as algebras $D$, $\widetilde{D}$, and $\overline{D}$ are isomorphic. Hence the map from $S_{m}^{G_{m}}/J$ to $S_{m}/I$ restricts to an isomorphism from $\widetilde{D}$ to $\overline{D}$ as $\mathcal{A}$-algebras. From the proof of Theorem 3.12, $\overline{D}$ is isomorphic to $(\mathbb{F}_{q}[W_{m}^{*}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}})^{q^{m-1}}$. Since $d_{1}(W_{m})$ generates $\mathbb{F}_{q}[W_{m}^{*}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}}$ (see [29]), $\overline{d_{1}}$ generates $\overline{D}$ and $\tilde{d_{1}}$ generates $\widetilde{D}$, as $\mathcal{A}$-algebras. From this we see that $\{\xi_{1},d_{1,m}\}$ generates $S_{m}^{G_{m}}$ as an $\mathcal{A}$-algebra. ∎
@@
$S_{m}^{G_{m}}$ is generated as an $\mathcal{A}$-algebra by $\{\xi_{1},\mathcal{R}(N(y_{1})^{q-1})\}$.
@@
Since $\deg(\mathcal{R}(N(y_{1})^{q-1}))=\deg(d_{1,m})<\deg(d_{i,m})$ for $i>1$, it follows from Theorem 4.8, that $\mathcal{R}(N(y_{1})^{q-1})=cd_{1,m}+\delta$ for some scalar $c$ and $\delta\in R_{n-1}$. By Theorem 5.23 and Lemma 3.11
@@ Acknowledgement
The computer algebra package Magma [2] was very useful in this work. It was used to test and confirm many hypotheses in low dimensions for small primes. We are grateful to Karl Dilcher and Keith Taylor (Canadian Mathematical Society book editors) and Donna Chernyk (Springer) for encouraging us to work on this problem. We thank the referee for a careful reading and constructive criticism.